# Align to Misalign: Automatic LLM Jailbreak with Meta-Optimized LLM Judges

### v1

@@ Abstract
Identifying the vulnerabilities of large language models (LLMs) is crucial for improving their safety by addressing inherent weaknesses. Jailbreaks, in which adversaries bypass safeguards with crafted input prompts, play a central role in red-teaming by probing LLMs to elicit unintended or unsafe behaviors. Recent optimization-based jailbreak approaches iteratively refine attack prompts by leveraging LLMs. However, they often rely heavily on either binary attack success rate (ASR) signals, which are sparse, or manually crafted scoring templates, which introduce human bias and uncertainty in the scoring outcomes. To address these limitations, we introduce AMIS (Align to MISalign), a meta-optimization framework that jointly evolves jailbreak prompts and scoring templates through a bi-level structure. In the inner loop, prompts are refined using fine-grained and dense feedback using a fixed scoring template. In the outer loop, the template is optimized using an ASR alignment score, gradually evolving to better reflect true attack outcomes across queries. This co-optimization process yields progressively stronger jailbreak prompts and more calibrated scoring signals. Evaluations on AdvBench and JBB-Behaviors demonstrate that AMIS achieves state-of-the-art performance, including 88.0% ASR on Claude-3.5-Haiku and 100.0% ASR on Claude-4-Sonnet, outperforming existing baselines by substantial margins.
@@ 3.1 Problem description
However, a manually constructed scoring template $\pi_{sc}$ may not generate a perfectly aligned score with the true ASR and could be suboptimal. Specifically, the score from the scoring template may be high, while the ASR of the prompt is still be 0. Thus, at the dataset level (Sec. 3.3), the goal of AMIS is to update and refine the scoring template itself so that it produces scores that are better aligned with the actual ASR results. The overall illustration and algorithm description are presented in Fig. 3 and Alg. 1, respectively.
@@ 4.1 Setups
We use Llama-3.1-8B-Inst. (Grattafiori et al., 2024) as the attacker model. For target models, we consider Llama-3.1-8B-Inst. (denoted as Llama-3.1-8B for convenience), GPT-4o-mini (gpt-4o-mini-2024-07-18) (OpenAI, 2024b), GPT-4o (gpt-4o-2024-11-20) (OpenAI, 2024a), Claude-3.5-Haiku (claude-3-5-haiku-20241022) (Anthropic, 2024), and Claude-Sonnet-4 (claude-sonnet-4-20250514) (Anthropic, 2025). The judge model is GPT-4o-mini in the inner loop, as well as for judging ASR. We use GPT-4o-mini for scoring template optimization in the outer loop. We initialize the framework with a prefix list of size $C=10$, and these are generated by GPT-4o (see Appendix I.5). Both the inner and outer loops are run for $L=5$ and $L^{\prime}=5$ iterations, respectively. In each inner loop, the attacker generates $M=5$ new candidate prompts, with the $K=5$ prompts provided as exemplar references to guide this generation. In each outer loop, the scoring template optimizer produces one new candidate template by conditioning on the entire history of previously used templates and their evaluation scores. We set the temperature of the attacker and the scoring template optimizer models to $1.0$ to encourage diverse generations, while all other models (targets, judges) use temperature $0.0$ for deterministic evaluation.
@@ 4.2 Main Results
Similarly, in JBB Behaviors (Table 2), our method maintains superior ASR while achieving notably higher StR scores, highlighting its ability to produce jailbreak prompts that are both more effective and more calibrated. Compared with the second-best method on this benchmark, it achieves an average gain of 20.2 in ASR and 0.41 in StR, further confirming the robustness and generality of our framework. These improvements are consistent across both open-weight (Llama-3.1-8B) and closed-source models (GPT-4o-mini, GPT-4o, Claude-3.5-Haiku, Claude-4-Sonnet), demonstrating the generalizability and transferability of our approach.
@@ 4.3 Additional analyses
We perform ablation studies to validate the proposed components of AMIS with the following five variants: (1) directly using $C$ initial prefixes (w/o inner, outer loop), (2) optimizing jailbreak prompts only with a fixed scoring template (w/o outer loop), (3) replacing our dense scoring template (Liu et al., 2024a) with the simpler ASR template while optimizing via the outer loop (w/o dense scoring template), (4) calculating alignment scores independently for each query without sharing templates (w/o dataset-level), and (5) not using prompt inheritance across iterations (w/o prompt inheritance). The experimental results are reported in Table 5. First, using only $C$ initial prefixes without any iterative refinement yields an ASR of 4.0 and a StR score of 0.04, revealing the necessity of optimization. While these values are very low, they indicate that the initial prefix set contains a small number of effective jailbreaks. Second, incorporating the outer loop to explicitly optimize the scoring template further increases performance, improving ASR by +2.0% and StR by +0.14. Third, replacing our dense scoring template with the simpler ASR template results in degraded performance (ASR: 74.0, StR: 0.40), suggesting that finer-grained rubrics provide more informative feedback to guide optimization. Fourth, replacing dataset-level optimization with query-level optimization leads to noticeably weaker results, underscoring the importance of leveraging cross-query signals rather than optimizing each query independently. Finally, removing prompt inheritance, i.e., restarting from the initial pool at every iteration, substantially reduces performance. Leveraging high-scoring prompts from previous iterations instead leads to significant improvements (ASR: 80.0 → 88.0, StR: 0.28 → 0.42). Together, these findings demonstrate that all components are essential to the effectiveness of our framework. Using only the initial prefixes without refinement yields very low performance (ASR: 4.0, StR: 0.04), showing the need for optimization. Optimizing the scoring template in the outer loop further improves both ASR and StR, while replacing it with a simpler ASR-only rubric degrades performance. Removing dataset-level optimization or prompt inheritance also leads to noticeably weaker results, underscoring the value of cross-query signals and iterative reuse of strong prompts. Together, these ablations confirm that each component contributes critically to the effectiveness of our framework.
@@
We next investigate the impact of the choice of attacker model. In addition to Llama-3.1-8B-Inst. used in our main experiments, we evaluate two stronger closed LLMs, GPT-4o-mini and Claude-3.5-Haiku, under the same experimental setting. The results are presented in Table 5. We find that GPT-4o-mini, which usually exhibits stronger safety alignment than Llama-3.1-8B-Inst., achieves lower ASR and StR scores. More surprisingly, Claude-3.5-Haiku shows even lower performance than Llama-3.1-8B-Inst. This is because highly safety-aligned models often refuse to produce harmful content during optimization, thereby limiting their effectiveness during the optimization to generate stronger jailbreak prompts on harmful queries.
@@
We further examine the role of the scoring template, which is a key component of our framework. To compare different designs, we run 25 ($L\times L^{\prime}$) iterations of jailbreak prompt optimization (in Sec. 3.2) using different variation of templates: ASR template ($\pi_{\mathrm{ASR}}$), StR template (Souly et al., 2024), and the AutoDAN-Turbo (AuT) template (Liu et al., 2024a), which serves as our initial template. The ASR template outputs binary scores (0 or 1), while the StR template provides scores in the range [0, 5], and the AuT template uses a [1,10] scale (prompt templates are provided in Appendix I). As shown in Table 5, the AuT template achieves the highest ASR and the second best StR. When further combined with outer-loop scoring template optimization (Sec. 3.3), the performance improves even more (see Table 5).
@@
As shown in Figure 5, prompts obtained from relatively less safety-aligned models, such as Llama and GPT series, exhibit poor transferability to more strongly safety-aligned models (Claude series). In contrast, prompts optimized on more strongly safety-aligned models like Claude-3.5-Haiku (Haiku-3.5) and Claude-4-Sonnet (Sonnet-4) transfer more effectively. Overall, these results suggest that prompts derived from more strongly safety-aligned models tend to generalize better across models, whereas those from weaker safety-aligned models remain highly model-specific. Interestingly, despite being the more recent model, Sonnet-4 not only achieves higher ASR in our main experiments but also shows lower transferability than Haiku-3.5. Paradoxically, Haiku-3.5 appears to demonstrate stronger safety alignment than Sonnet-4 in our experiments, highlighting that model updates do not always lead to consistent improvements in robustness against jailbreak transferability. Additional results on Sonnet-3.5 are presented in Appendix C.
@@ 5 Conclusion
While AMIS achieves consistent improvements across diverse benchmarks, several limitations remain. First, our evaluation of ASR relies on an LLM-as-a-judge setup following the prior work, which may introduce inherent biases and limit the robustness. Second, the iterative inner–outer loop optimization entails non-trivial computational costs. However, we note that our approach is even cheaper than the baseline approach in achieving the same ASR (see Appendix E for detailed experiments). We expect future works to mitigate these limitations by incorporating multi-judge or human-in-the-loop evaluation and developing more sample-efficient optimization strategies.
@@
We provide detailed implementation information, including prompt designs, APIs, and hyperparameter settings, as well as experimental setups such as datasets and evaluation metrics in Section 4 and Appendix I. Furthermore, we will release the source code in the near future.
@@ Appendix A Algorithmic Details of AMIS
In this section, we present algorithm of AMIS in Alg. 1. AMIS is a meta-optimization framework that alternates between two levels of optimization. At the query level (inner loop), we iteratively refine jailbreak prompts for each harmful query: an attacker LLM generates candidate prompts, the target model produces responses, and a judge model scores them with a fine-grained scoring template, retaining only the top-ranked candidates across iterations. At the dataset level (outer loop), we update the scoring template itself so that its scores better align with true binary ASR outcomes. This meta-optimization process ensures that both the prompts and the evaluation rubric co-evolve, yielding stronger jailbreak attacks and more reliable optimization signals.
@@ B.2 Evaluation with HarmBench ASR prompt
Since our approach relies on the ASR alignment score, it employs the ASR prompt internally, which is also used for evaluation. To ensure a fair comparison, we additionally adopt alternative ASR prompt at the evaluation, which is not seen during the optimization. Specifically, we adopt an ASR prompt from HarmBench [^2] https://github.com/centerforaisafety/HarmBench/blob/main/eval_utils.py while using the same GPT-4o-mini as the judge model. The results on AdvBench and JBB-Behaviors are reported in Tables 7 and 8, respectively. Across all five models, our method consistently achieves the highest ASR, demonstrating its robustness and effectiveness.
@@
Attacks $\downarrow$ / Targets $\rightarrow$ Llama-3.1-8B GPT-4o-mini GPT-4o Haiku-3.5 Sonnet-4 Vanilla 36.0 4.0 0.0 0.0 0.0 PAIR 88.0 74.0 62.0 42.0 22.0 TAP 96.0 80.0 46.0 32.0 24.0 PAP 56.0 44.0 44.0 10.0 10.0 SeqAR 90.0 42.0 0.0 0.0 6.0 AutoDAN-Turbo 88.0 56.0 40.0 42.0 42.0 AMIS (Ours) 98.0 98.0 98.0 72.0 94.0
@@
Attacks $\downarrow$ / Targets $\rightarrow$ Llama-3.1-8B GPT-4o-mini GPT-4o Haiku-3.5 Sonnet-4 Vanilla 47.0 11.0 7.0 4.0 5.0 PAIR 92.0 78.0 70.0 52.0 30.0 TAP 93.0 81.0 65.0 45.0 23.0 PAP 99.0 87.0 83.0 74.0 32.0 SeqAR 93.0 0.0 0.0 3.0 14.0 AutoDAN-Turbo 86.0 61.0 51.0 35.0 36.0 AMIS (Ours) 100.0 99.0 93.0 78.0 88.0
B.3 Case studies on prefixes
As described in Sections 4.1 and I.5, we use 10 prefixes generated by GPT-4o, to concatenate with the query as the initial prompts. To examine the effectiveness of these prefixes, we adopt the “Human” baseline from Mazeika et al. (2024), which employs a fixed set of in-the-wild human jailbreak templates similar to the Do Anything Now (DAN) jailbreaks. For both Case 1 and Case 2, we randomly sample 10 prefixes without overlap between the two cases. The results are summarized in Table 9. In Case 1, the model with prefixes alone fails to produce any successful attacks (ASR = 0.0, StR = 0.0), whereas applying AMIS increases the ASR and StR to 52.0 and 0.26, respectively. In Case 2, the prefixes alone already yield relatively strong performance (ASR = 74.0, StR = 0.16), and the use of AMIS further improves them to 90.0 and 0.50. These results indicate that AMIS consistently enhances both ASR and StR regardless of the quality (i.e., those that can already achieve higher ASR values) of the initial prefixes. Nevertheless, prefixes with higher quality appear to contribute more effectively to overall performance gains.
Table 9: Case study on the effectiveness of prefixes
ASR StR

### v2

@@ Abstract
Identifying the vulnerabilities of large language models (LLMs) is crucial for improving their safety by addressing inherent weaknesses. Jailbreaks, in which adversaries bypass safeguards with crafted input prompts, play a central role in red-teaming by probing LLMs to elicit unintended or unsafe behaviors. Recent optimization-based jailbreak approaches iteratively refine attack prompts by leveraging LLMs. However, they often rely heavily on either binary attack success rate (ASR) signals, which are sparse, or manually crafted scoring templates, which introduce human bias and uncertainty in the scoring outcomes. To address these limitations, we introduce AMIS (Align to MISalign), a meta-optimization framework that jointly evolves jailbreak prompts and scoring templates through a bi-level structure. In the inner loop, prompts are refined using fine-grained and dense feedback from a fixed scoring template. In the outer loop, the template is optimized using an ASR alignment score, gradually evolving to better reflect true attack outcomes across queries. This co-optimization process yields progressively stronger jailbreak prompts and more calibrated scoring signals. Evaluations on AdvBench and JBB-Behaviors demonstrate that AMIS achieves state-of-the-art performance, including 88.0% ASR on Claude-3.5-Haiku and 100.0% ASR on Claude-4-Sonnet, outperforming existing baselines by substantial margins. [^1] Code available at: https://github.com/hamin2065/AMIS
@@ 3.1 Problem description
However, a manually constructed scoring template $\pi_{sc}$ may not generate a perfectly aligned score with the true ASR and could be suboptimal. Specifically, the score from the scoring template may be high, while the ASR of the prompt is still 0. Thus, at the dataset level (Sec. 3.3), the goal of AMIS is to update and refine the scoring template itself so that it produces scores that are better aligned with the actual ASR results. The overall illustration and algorithm description are presented in Fig. 3 and Alg. 1, respectively.
@@ 4.1 Setups
We use Llama-3.1-8B-Inst. (Grattafiori et al., 2024) as the attacker model. For target models, we consider Llama-3.1-8B-Inst. (denoted as Llama-3.1-8B for convenience), GPT-4o-mini (gpt-4o-mini-2024-07-18) (OpenAI, 2024a), GPT-4o (gpt-4o-2024-11-20) (OpenAI, 2024b), Claude-3.5-Haiku (claude-3-5-haiku-20241022) (Anthropic, 2024), and Claude-Sonnet-4 (claude-sonnet-4-20250514) (Anthropic, 2025). The judge model is GPT-4o-mini in the inner loop, as well as for judging ASR. We use GPT-4o-mini for scoring template optimization in the outer loop. We initialize the framework with a prefix list of size $C=10$, and these are generated by GPT-4o (see Appendix I.5). Both the inner and outer loops are run for $L=5$ and $L^{\prime}=5$ iterations, respectively. In each inner loop, the attacker generates $M=5$ new candidate prompts, with the $K=5$ prompts provided as exemplar references to guide this generation. In each outer loop, the scoring template optimizer produces one new candidate template by conditioning on the entire history of previously used templates and their evaluation scores. We set the temperature of the attacker and the scoring template optimizer models to $1.0$ to encourage diverse generations, while all other models (targets, judges) use temperature $0.0$ for deterministic evaluation.
@@ 4.2 Main results
Similarly, in JBB Behaviors (Table 2), our method maintains superior ASR while achieving notably higher StR scores, highlighting its ability to produce jailbreak prompts that are both more effective and more calibrated. Compared with the second-best method on this benchmark, it achieves an average gain of 20.2% in ASR and 0.41 in StR, further confirming the robustness and generality of our framework. These improvements are consistent across both open-weight (Llama-3.1-8B) and closed-source models (GPT-4o-mini, GPT-4o, Claude-3.5-Haiku, Claude-4-Sonnet), demonstrating the generalizability and transferability of our approach.
@@ 4.3 Additional analyses
We perform ablation studies to validate the proposed components of AMIS with the following five variants: (1) directly using $C$ initial prefixes (w/o inner, outer loop), (2) optimizing jailbreak prompts only with a fixed scoring template (w/o outer loop), (3) replacing our dense scoring template (Liu et al., 2024a) with the simpler ASR template while optimizing via the outer loop (w/o dense scoring template), (4) calculating alignment scores independently for each query without sharing templates (w/o dataset-level), and (5) not using prompt inheritance across iterations (w/o prompt inheritance). The experimental results are reported in Table 5. First, using only $C$ initial prefixes without any iterative refinement yields an ASR of 4.0 and a StR score of 0.04, revealing the necessity of optimization. While these values are very low, they indicate that the initial prefix set contains a small number of effective jailbreaks. Second, incorporating the outer loop to explicitly optimize the scoring template further increases performance, improving ASR by +2.0% and StR by +0.14. Third, replacing our dense scoring template with the simpler ASR template results in degraded performance (ASR: 74.0, StR: 0.40), suggesting that finer-grained rubrics provide more informative feedback to guide optimization. Fourth, replacing dataset-level optimization with query-level optimization leads to noticeably weaker results, underscoring the importance of leveraging cross-query signals rather than optimizing each query independently. Finally, removing prompt inheritance, i.e., restarting from the initial pool at every outer iteration, substantially reduces performance. Leveraging high-scoring prompts from previous iterations instead leads to significant improvements (ASR: 80.0 → 88.0, StR: 0.28 → 0.42). Together, these findings demonstrate that all components are essential to the effectiveness of our framework. Using only the initial prefixes without refinement yields very low performance (ASR: 4.0, StR: 0.04), showing the need for optimization. Optimizing the scoring template in the outer loop further improves both ASR and StR, while replacing it with a simpler ASR-only rubric degrades performance. Removing dataset-level optimization or prompt inheritance also leads to noticeably weaker results, underscoring the value of cross-query signals and iterative reuse of strong prompts. Together, these ablations confirm that each component contributes critically to the effectiveness of our framework.
@@
We next investigate the impact of the choice of attacker model. In addition to Llama-3.1-8B used in our main experiments, we evaluated two stronger closed LLMs, GPT-4o-mini and Claude-3.5-Haiku, under the same experimental setting. The results are presented in Table 5. We find that GPT-4o-mini, which usually exhibits stronger safety alignment than Llama-3.1-8B, achieves lower ASR and StR scores. More surprisingly, Claude-3.5-Haiku shows even lower performance than Llama-3.1-8B. This is because highly safety-aligned models often refuse to produce harmful content during optimization, thereby limiting their effectiveness during the optimization to generate stronger jailbreak prompts on harmful queries.
@@
We further examine the role of the scoring template, which is a key component of our framework. To compare different designs, we ran 25 ($L\times L^{\prime}$) iterations of jailbreak prompt optimization (in Sec. 3.2) using different variations of templates: ASR template ($\pi_{\mathrm{ASR}}$), StR template (Souly et al., 2024), and the AutoDAN-Turbo (AuT) template (Liu et al., 2024a), which serves as our initial template. The ASR template outputs binary scores (0 or 1), while the StR template provides scores in the range [0, 5], and the AuT template uses a [1,10] scale (prompt templates are provided in Appendix I). Among these three templates, the AuT template achieves the highest ASR and StR, as shown in Table 5. When further combined with outer-loop scoring template optimization (Sec. 3.3), the performance improves even more (see Table 5).
@@
As shown in Figure 5, prompts obtained from relatively less safety-aligned models, such as Llama and GPT series, exhibited poor transferability to more strongly safety-aligned models (Claude series). In contrast, prompts optimized on more strongly safety-aligned models like Claude-3.5-Haiku (Haiku-3.5) and Claude-4-Sonnet (Sonnet-4) transferred more effectively. Overall, these results suggest that prompts derived from more strongly safety-aligned models tend to generalize better across models, whereas those from weaker safety-aligned models remain highly model-specific. Interestingly, despite being the more recent model, Sonnet-4 not only achieved higher ASR in our main experiments but also showed lower transferability than Haiku-3.5. Paradoxically, Haiku-3.5 appears to demonstrate stronger safety alignment than Sonnet-4 in our experiments, highlighting that model updates do not always lead to consistent improvements in robustness against jailbreak transferability. Additional results on Sonnet-3.5 are in Appendix C.
Cross-dataset generalization.
To assess whether AMIS overfits to a specific dataset, we run cross-dataset transfer experiments using AdvBench and JBB-Behaviors (JBB). In this setup, scoring templates optimized on one dataset are applied directly to the other without any further adaptation. We observe strong transfer in both directions. Templates trained on JBB reach an ASR of 86% on AdvBench compared with 88% for the original setting. StR remains stable across settings. Interestingly, we find that AdvBench-optimized template performs even better on JBB, reaching an ASR of 87% compared with the original JBB score of 78%. We conjecture this is due to the different nature of the two benchmarks. AdvBench contains short and direct harmful instructions across 31 categories, which provide clean signals that help the template learn broad and well-calibrated harmfulness criteria. JBB includes nearly 100 categories, covering social engineering, deception, misinformation, and physical harm, which creates a more heterogeneous distribution. The clearer structure of AdvBench therefore produces a template with generalizable cues that transfer effectively to the more diverse JBB setting. These findings indicate that AMIS learns generalizable harmfulness criteria rather than memorizing dataset-specific patterns, demonstrating robust cross-dataset generalization.
Table 6: Cross-dataset generalization.
ASR StR
JBB $\rightarrow$ AdvBench 86.0 0.42
AdvBench (original) 88.0 0.42
AdvBench $\rightarrow$ JBB 87.0 0.50
JBB (original) 78.0 0.48
Figure 6: Cost analysis: PAIR vs AMIS
Cost analysis.
We conduct a cost comparison using Claude-4-Sonnet as the target model on AdvBench. To ensure a fair comparison, we extend the number of iterations for PAIR. As shown in Figure 6, although there is a small overlap at the very beginning, AMIS consistently exhibits a much higher ASR relative to cost throughout the iterations. AMIS reaches 100% ASR at approximately \$1.96 per query, whereas PAIR achieves only 94% ASR even at a higher cost. This result demonstrates that even though AMIS requires a comparable amount of total cost, it achieves full coverage with faster ASR growth and better cost-efficiency. Overall, this experiment indicates that AMIS provides a more efficient optimization trajectory than PAIR under similar computational budgets.
(a) ASR alignment score across outer iterations. Trajectories over the outer loop illustrating alignment dynamics for (a) Llama-3.1-8B, (b) Claude-3.5-Haiku, and (c) Claude-4-Sonnet.
Convergence of ASR alignment score.
To analyze how scoring templates evolve during outer-loop meta-optimization, we report the ASR Alignment Score at each iteration, separately for ASR=1 (successful jailbreaks) and ASR=0 (refusal cases). Figure 9(a) shows the trajectories for three models on AdvBench over 10 outer iterations. Across all models, ASR=1 scores increase steadily, indicating that meta-optimization progressively refines the templates to assign higher alignment scores to harmful responses. In contrast, ASR=0 scores remain largely stable, with only a small shift at the first iteration corresponding to the transition from the handcrafted template to the first attacker-aware version. Subsequent iterations mainly introduce incremental refinements, leading to stable ASR=0 behavior. Overall, the rising ASR=1 scores alongside stable ASR=0 scores suggest that optimization improves sensitivity to harmful content while maintaining consistency for refusal cases. Additional analyses are provided in the Appendix E.
@@ 5 Conclusion
While AMIS achieves consistent improvements across diverse benchmarks, several limitations remain. First, our evaluation of ASR relies on an LLM-as-a-judge setup following the prior work, which may introduce inherent biases and limit the robustness. Second, the iterative inner–outer loop optimization entails non-trivial computational costs. However, we note that our approach is even cheaper than the baseline approach in achieving the same ASR (Figure 6). We expect future works to mitigate these limitations by incorporating multi-judge or human-in-the-loop evaluation and developing more sample-efficient optimization strategies.
@@
We provide detailed implementation information, including prompt designs, APIs, and hyperparameter settings, as well as experimental setups such as datasets and evaluation metrics in Section 4 and Appendix I. The source code of AMIS is publicly available at https://github.com/hamin2065/AMIS.
Acknowledgments
Both Hamin Koo and Jaehyung Kim are affiliated with the Department of Artificial Intelligence at Yonsei University. This research was supported in part by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University); No. RS-2025-02215344, Development of AI Technology with Robust and Flexible Resilience Against Risk Factors).
@@ Appendix A Algorithmic Details of AMIS
AMIS is a meta-optimization framework that alternates between two levels of optimization. At the query level (inner loop), we iteratively refine jailbreak prompts for each harmful query: an attacker LLM generates candidate prompts, the target model produces responses, and a judge model scores them with a fine-grained scoring template, retaining only the top-ranked candidates across iterations. At the dataset level (outer loop), we update the scoring template itself so that its scores better align with true binary ASR outcomes. This meta-optimization process ensures that both the prompts and the evaluation rubric co-evolve, yielding stronger jailbreak attacks and more reliable optimization signals.
@@ Appendix B Additional Experiments
We conduct additional analyses to expand our understanding of AMIS. Results are based on AdvBench with Claude-3.5-Haiku unless otherwise specified and are reported using attack success rate (ASR) and strongREJECT score (StR).
@@ B.2 Evaluation with HarmBench ASR Prompt
Since our approach relies on the ASR alignment score, it employs the ASR prompt internally, which is also used for evaluation. To ensure a fair comparison, we additionally adopt alternative ASR prompt that is not seen during the optimization process. Specifcially, we adopt ASR prompt from HarmBench [^3] https://github.com/centerforaisafety/HarmBench/blob/main/eval_utils.py while using the same GPT-4o-mini as the judge model. The results on AdvBench and JBB-Behaviors are reported in Tables 8 and 9, respectively. Across all five models, our method consistently achieves the highest ASR, demonstrating its robustness and effectiveness.
@@
Attacks $\downarrow$ / Targets $\rightarrow$ Llama-3.1-8B GPT-4o-mini GPT-4o Haiku-3.5 Sonnet-4 Vanilla 36.0 4.0 0.0 0.0 0.0 PAIR 88.0 74.0 62.0 42.0 22.0 TAP 96.0 80.0 46.0 32.0 24.0 PAP 56.0 44.0 44.0 10.0 10.0 SeqAR 90.0 42.0 0.0 0.0 6.0 AutoDAN-Turbo 88.0 56.0 40.0 42.0 42.0 AMIS 98.0 98.0 98.0 72.0 94.0
@@
Attacks $\downarrow$ / Targets $\rightarrow$ Llama-3.1-8B GPT-4o-mini GPT-4o Haiku-3.5 Sonnet-4 Vanilla 47.0 11.0 7.0 4.0 5.0 PAIR 92.0 78.0 70.0 52.0 30.0 TAP 93.0 81.0 65.0 45.0 23.0 PAP 99.0 87.0 83.0 74.0 32.0 SeqAR 93.0 0.0 0.0 3.0 14.0 AutoDAN-Turbo 86.0 61.0 51.0 35.0 36.0 AMIS 100.0 99.0 93.0 78.0 88.0
B.3 Additional Experiments on Optimization Models
Table 10: Performance of AMIS under different optimization model configurations on AdvBench. ASR and StR are reported for each setup.